OpenStudy (anonymous):

for what value of p and q will the following pairs of linear equations has infinitely many solutions? 4x+5y=2;(2p+7q)x+(p+8q)y=2q-p+1

OpenStudy (asnaseer):

what are your thoughts on how to do this?

OpenStudy (anonymous):

on how to do it

OpenStudy (asnaseer):

do you know under what conditions a pair of linear equations has infinitely many solutions?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

could you explain please?

OpenStudy (asnaseer):

(I am trying to guide you to how to solve this question)

OpenStudy (anonymous):

a1/a2=b1/b2=c1/ c2

OpenStudy (asnaseer):

perfect! :)

OpenStudy (asnaseer):

so if the coefficients of the second equation are a multiple of the coefficients of the first equation, then the two equations are not linearly independent and will therefore have an infite number of solutions.

OpenStudy (asnaseer):

so you can start by saying that all the coefficients of the second equation must some multiple of the first equation - lets cal this multiple 'k' this gives us the following 3 equations to solve:\[\begin{align} 2p+7q&=4k\tag{1}\\ p+8q&=5k\tag{2}\\ 2q-p+1&=2k\tag{3}\end{align}\]make sense?

OpenStudy (anonymous):

yah got it

OpenStudy (asnaseer):

hopefully you can solve from here?

OpenStudy (anonymous):

thanks a lot

OpenStudy (asnaseer):

yw :)